Department of Computer Science

University of Edinburgh

Mission Statement

The University of Edinburgh is internationally recognised as a leading research university. It is building its future on strength in research, and on the promotion of education and training whose qualities depend on a challenging environment of research and professional practice that results in a distinctive type of teaching. Within the University, the Faculty of Science and Engineering is a particular advocate of the benefits of teaching in a research environment.

The Department of Computer Science, together with the Department of Artificial Intelligence, the Centre for Cognitive Science and the Human Communication Research Centre, form the Informatics Planning Unit within the Faculty of Science and Engineering. The emerging subject of informatics is a discipline that involves the study of the processes underlying information, its acquisition, communication and manipulation, and the tools needed for these processes. At Edinburgh, computer science research has been mostly concerned with engineering principles and the rigorous mathematisation of concepts that lie behind methods and languages for hardware and software design. This has complemented work carried out by the other members of the Planning Unit.

Already at the leading edge in many separate informatics topics, Edinburgh has the potential to be a world leader in making informatics an integrated science and engineering discipline in a broad sense: foundational, engineering, experimental, cognitive and social. Such a goal will ensure that applications develop with a foundation of coherent concepts.

The Informatics Planning Unit aims to share organisational mechanisms where appropriate, encourage research collaboration within the unit, establish a graduate school and develop an integrated teaching programme. However, the identity of separate departments and groupings will be retained as appropriate to their different contributions to informatics.

In the Department of Computer Science, the mutual dependence between scientific theory and engineering practice is an important underpinning principle. More specifically,

• computer science is the study of the structure and dynamics of computational systems using mathematical, logical and empirical tools;

• engineering is the application of scientific understanding and analysis to the design and implementation of useful systems; and

• computer science must be informed by, and should cast light upon, the engineering design problems relating to the exploitation of computational systems in the human environment.

This viewpoint informs both research and education activities.

The Department of Computer Science intends to remain at the forefront of research in its areas of expertise, primarily in theoretical computer science organised within the Laboratory for Foundations of Computer Science and in computer systems organised within the Computer Systems Group. The Department also encourages and supports individuals whose work falls outside these main groupings. A European dimension to research, and research links with industry, are both being developed strongly.

In teaching, the Department aims to develop well-rounded graduates with a high level of knowledge and skills, both practical and theoretical, by ensuring that students benefit from its strengths as a leading research department. In addition to technical skills, the Department aims to develop graduates with appropriate general skills and attitudes expected of a professional scientist or engineer. The undergraduate curriculum is designed to provide a progressive development of study skills, conceptual framework, and practical skills and experience.